OpenStudy (anonymous):

What belongs in the box marked with the question mark in the proof? Given: 3x - 4 = 2(x + 8), Prove: x = 20 (1) 3x - 4 = 2(x + 8) Blank (2) 3x - 4 = 2x + 16 Blank (3) 3x = 2x + 20 Blank (4) x = 20 ?

OpenStudy (anonymous):

the blank boxes are for you to plug in x=20 to prove its right. so it would be 3(20)-4=2(20+8) 60-4=2(28) 56=56 so its true!

OpenStudy (anonymous):

i need the reason =S

OpenStudy (anonymous):

you need the reason? i just proved that it was correct?

OpenStudy (anonymous):

http://prntscr.com/cojcl that is a similar image to what i need i neet the answer to the one under the question mark in the photo

OpenStudy (anonymous):

it has to be one of the propertyies: substitution adition multiplication etc... and i dont know what it is

OpenStudy (anonymous):

im pretty sure its the associative property of multiplication